OpenStudy (anonymous):

A train one mile long travels at 30 mph. It enters a tunnel one mile long at 4pm. At what time does the end of the train emerge from the tunnel.

OpenStudy (anonymous):

r*t = d The rear end of the train travels 1 mile to enter the 1 mile long tunnel and another mile to exit the tunnel. 30*t=2 t=2/30 hour = 4/60 hour = 4 minutes. The clock will indicate 4:04 PM when the train emerges from the tunnel.

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

Your welcome and thanks for the answer.